OpenStudy (anonymous):

determine whether the following side measures form right angles. 4,7, radical 65

myininaya (myininaya):

which is the biggest number?

myininaya (myininaya):

is that sqrt(65)?

myininaya (myininaya):

i assume you mean sqrt(65)

myininaya (myininaya):

since sqrt(65) is the biggest, let's assume thats our hypothnues use the pytagorean thm

myininaya (myininaya):

does 4^2+7^2=(sqrt(65))^2 16+49=65? 65=65 so it is a right triangle

OpenStudy (anonymous):

ahh yeah! haha thanks so much